# Combining Threat Intelligence with IoT Scanning to Predict Cyber Attacks

### v3

@@ Abstract
keywords:
Internet of Things, Dark Web, IoT Vulnerability, Machine Learning, Cybersecurity, Threat Intelligence
1 Introduction
@@
In this paper, I focus on the Internet of Things (IoT) as a vulnerability due to its expanding presence and susceptibility to exploitation [2]. By employing the structured methodology outlined in this paper, I evaluate the degree to which cybercriminal groups target IoT systems. Through my analysis of 23 criminal websites on the Dark Web, I discovered that even seemingly harmless devices, such as small sensors, can cause substantial damage when compromised. IoT opens up an entirely new realm for cybercriminals to exploit. These hacker groups generally do not discriminate in their targets—they exploit any weakness or vulnerability they can find.
@@
The remainder of the paper is organized as follows: the second section provides an overview of the concepts introduced in this paper; the third section outlines a methodology for collecting, analyzing Dark Web data, and identifying vulnerabilities[3]. The fourth section details how this methodology was applied in a case study of IoT using Shodan, the IoT search engine. The final section concludes the study and explores potential solutions.
2 Literature Review
2.1 Dark Web Utilization
@@
Hackers, criminals, and terrorists leverage the web to facilitate their illicit activities. Since the emergence of the Dark Web, numerous new marketplaces have emerged, offering a wide and often alarming array of products, including prescription pills, meth, heroin, speed, crack, guns, stolen identities, gold, and erotica. The Dark Web is also a platform for spreading extremism, facilitating communication among terrorists and criminals, and sharing knowledge about illegal hacks and attacks, thereby enabling harmful activities that threaten society. Most of these markets operate using Bitcoin and other cryptocurrencies that allows for anonymous online transactions[4]. According to one study, criminals and terrorists use Bitcoin to sell drugs and subsequently purchase guns and ammunition. During my exploration of the dark market known as Hansa Market (http://hansamkt2rr6nfg3.onion/search/?q=hacker&c=59), I identified a website called Arms and Ammunition (http://armsforsd.com/index.html), which poses significant threats to nations both internally and externally. This underscores the importance of developing robust threat intelligence.
2.2 Discovering IoT Devices Through Shodan
I chose to focus on IoT devices due to their profound impact on various aspects of daily life. These devices, similar to our mobile phones, will become integral to our routines and will predict our behavior. Despite their numerous applications, IoT devices are highly susceptible to hacks and malfunctions. Many of these devices operate on low-power CPUs and custom operating systems that lack the capability to support encryption and decryption mechanisms, making them inherently insecure. For instance, a smart thermostat does more than just regulate temperature; it stores data on a central server that helps predict user temperature patterns. While a compromised thermostat in a home may not be immediately dangerous, in industrial control systems, even minor data manipulation through sensors can have catastrophic consequences. If hackers or criminals exploit the vulnerabilities of these critical IoT devices, they could potentially threaten individuals based on their past behavior or disrupt essential services like power stations[5]. To identify vulnerable IoT devices, I utilized Shodan. Shodan is an IoT search engine that scans the internet to locate devices with open ports and establish successful connections. It can identify a wide range of internet-connected devices, including computer servers, webcams, industrial control systems, power plants, traffic signals, and virtually any other device. Shodan is a relatively new search engine specifically designed for IoT devices, and its reliability has been validated by several research studies. One such study, conducted over several months, connected a few devices to the internet for the first time to see if Shodan would list them. After 14 days, the study found that the devices’ information, including IP addresses, ports, locations, and device details, appeared on Shodan [6]. As per findings, Shodan collects this information through service banner interrogation, which involves extracting data from a system in response to a connection request. Banner grabbing is the process of obtaining active information about a service or system through port interrogation. With various filters available - such as IP address, port, country, city, latitude, longitude, and device information (OS) - one can identify and exploit vulnerable devices[7]. However, this capability is invaluable for developing threat intelligence.
2.3 Developing Cyber Threat Intelligence
@@
My research focused on determining whether the systematic collection and processing of content from the Dark Web could contribute to the development of actionable threat intelligence for protecting vulnerable IoT devices. I specifically examined how insights gathered from underground hacker forums, marketplaces, and other illicit platforms could enhance the understanding of emerging cyber threats targeting IoT ecosystems[8, 9]. The remainder of this paper details the methodologies I employed for data collection, the techniques I used to filter and refine relevant information, and the analytical approaches I applied to transform raw data into meaningful threat intelligence. By outlining these processes, this study demonstrates how Dark Web intelligence could have been leveraged to strengthen IoT security measures and mitigate cyber risks before they materialized into active threats.
3 Methodology
3.1 Data Collection Framework
@@
To address the diverse threats posed by the information sources that hackers, criminals, and terrorists use to conduct destructive activities, I propose a comprehensive threat intelligence framework. This framework integrates Dark Web data aggregation, machine learning-driven anomaly detection for IoT vulnerability patterns, and human-AI collaboration for real-time cyber-attack prediction.
@@
Unlike previous methodologies [10], this approach emphasizes the correlation of hacker discussions (e.g., IoT exploitation trends) with Shodan-derived device vulnerabilities to prioritize high-risk targets. Figure 1 illustrates the methodology, which is designed to effectively assist human investigators in obtaining Dark Web intelligence through various information sources, collection methods, filtering, and analysis [11].
  • Information Sources encompass a broad spectrum of providers of hacker-related information on the web. These sources range from readily accessible platforms, such as search engines, to websites developed and maintained by hacker groups, hacktivists, terrorist organizations[12], or their supporters. Access to the latter often requires the assistance of domain experts.
  • Collection methods enable the systematic searching and browsing of information from identified sources[13]. IoT Threat Horizon Scraping begins with a set of relevant seed URLs or keywords, which are derived from the initial information gathered through browsing. IoT threat scraping can be done using various web crawling tools.
  • Back-link search, supported by search engines such as Google (www.google.com) and AltaVista (www.altavista.com, which was acquired by Yahoo! in 2003), allows investigators to find web pages that link to a target web domain or page. This feature is particularly useful for tracing the activities of hacker supporters and sympathizers, whose web pages often reference hacker sites[14]. For example, these pages may glorify hacking actions or demonstrate a pattern of coordinated attacks.
  • Personal/People Data Search, exemplified by major web portals such as Yahoo! (members.yahoo.com) and MSN (groups.msn.com), uncovers the profiles of groups or individuals with shared interests. Hackers and their supporters may include ”hot links” in their profiles[15], which enable investigators to uncover hidden connections.
  • Meta-searching leverages related keywords as input to query multiple search engines, allowing investigators or automated programs to collate top-ranked results and eliminate duplicates to obtain highly relevant information. For example, searching for the hacker group ”Anonymous” (widely known and famous for attacking the FBI, CIA, PayPal, Sony, MasterCard, Visa, Chinese, and Israeli governments) across multiple search engines yields mixed results about their hacking news articles and websites. Augmenting ”Anonymous” with terms like ”digital Robin Hood,” ”freedom fighters,” or ”Operation Anons” (as glorified by their supporters) can yield more targeted and relevant results[16].
  • Filtering involves sorting through collected information and eliminating irrelevant results. This task requires domain knowledge and linguistic expertise[17]. Domain knowledge encompasses understanding hackers, terrorists, and their supporter groups, including their presence on and usage of the web, as well as their histories, activities, and missions. Linguistic knowledge pertains to recognizing terms, slogans, and other textual and symbolic clues in the native language of cyber-attackers.
  • Analysis provides insights into data and aids investigators in identifying trends and verifying conjectures. Several functions support these analytical tasks. IoT Exploit Motive Profiling identifies common properties among entities and categorizes them to help investigators predict trends[18]. Clustering organizes entities into naturally occurring groups, facilitating the identification of similar hacker groups. IoT Risk Topography Mapping presents data in a visually comprehensible format, enabling investigators to visualize relationships within the networked structure of these hacker groups or their activities. Threat Scoring ranks vulnerabilities based on hacker interest and entity exposure. Vulnerability Discovery presents methods for selecting an entity and uncovering various threats that could be exploited for that chosen entity.
3.2 Implementation Approach
@@
As discussed earlier, the proposed methodology integrates a combination of advanced data mining, web mining, and machine learning techniques while leveraging human domain expertise to refine their application. This hybrid approach enhances the ability to systematically identify and monitor criminal groups, analyze their discussions, and extract intelligence relevant to emerging cyber threats[20]. By incorporating both automated data processing and human analytical oversight, this methodology balances computational efficiency with contextual accuracy, enabling more precise intelligence gathering[4]. Ultimately, this semi-automated framework presents a scalable solution for uncovering hidden threat actors and understanding the evolving landscape of cybercrime on the Dark Web.
4 Case Study: IoT Scanning to Predict Cyber Attacks
To illustrate the effectiveness of the proposed methodology, we applied it to collect and analyze web-based activities of hackers, terrorists, and criminals targeting IoT, a technology with inherent security vulnerabilities. Initially, we identified dark-market sites, forums, and blogs, followed by an analysis of IoT exploitation through these platforms. Research suggests that nearly 57% of IoT devices are vulnerable [17], despite representing just 0.1% of all internet-connected devices. By 2030, projections indicate that over 26 billion IoT devices will be online, expanding the attack surface for cybercriminals. Given these trends, our case study focused on IoT security threats.
4.1 Collection
@@
To gather data, I first identified several suspicious URLs using web searches, referencing dark web markets[21, 22]. I conducted web spidering using the ’Open Web Spidering’ tool. For instance, starting with keywords like ’freedom fighters’ and ’digital Robin Hood,’ I found four intriguing URLs: HackHound (a forum for hackers), Du Bist Anonymous (a website supporting the Anonymous group), CiHad.net (a suspicious website associated with a terror group), and Intel Exchange (a political ideology website). Using these categories as references, we uncovered additional websites during data collection. I then performed the same search on dark web search engines such as Torch, Grams, and Ahmia, although the results were scattered. I stored them nonetheless, as they primarily pointed to dark web markets. Manually browsing these markets revealed offers of both hacking services and services required for hacking. I selected keywords such as ’hacking services,’ ’free world,’ ’personal army,’ and ’Internet of Things,’ yielding various results. Overall, I collected information from hundreds of sources, and it would be an interesting future direction to study how data collection techniques affect the quality of analysis.
4.2 Filtering
@@
I conducted two rounds of filtering. First, I manually removed unrelated sites, such as news, games, and porn sites, which had no reference to hacking. I retained websites of forums, blogs, ideological discussions, dark markets for hacking and ammunition services. After this round, thirty-two sites remained. Second, given my limited domain expertise and knowledge of languages, I initially manually removed some websites that felt less relevant and similar to English forum websites in my results. Subsequently, I used natural language processing (NLP) and machine learning (ML) for keyword tagging and threat scoring.
4.3 Analysis
4.3.1 Classification and Usage of Websites
@@
I performed clustering and classification of the remaining 23 websites. These websites provided a diverse range of information necessary for my analysis. Table 1 lists these websites with hyperlinked URLs and descriptions.
@@
Table 1: Websites collected with the implementation of the methodology.
I followed following steps to analyze “IoT Hacking” or “hacking IoT devices” discussions on some of these websites: Filtered Cyber Security Hacking Discussion Forums from above list ignoring rest of the websites from my list. Manually browsed through these forums and searched for above keywords in these forums. Gathered the statistics of as recent (to 2024) as possible and generated a graph, shown in Figure 2, and a table shown in Table 2. The graph in Figure 2 displays the percentage of IoT hacking-related discussions occurring in the seven forums I chose for analysis. The x-axis shows the names of these forums, while the y-axis shows the total percentage of IoT hacking discussions. The ”Hacker Web” forum, which is a collection of 18 major forums, has 3.3
@@
Table 2: This table shows the keywords searched and data found on the mentioned websites.
By manually browsing through these forums, I analyzed the most frequently discussed topics related to IoT hacking. I selected four websites listed in Table 2, searched for the keywords in the ’Searched’ column, and recorded the results and total post counts in the ’Found’ and ’# of Posts’ columns, respectively. My analysis revealed that botnets, malware, and sensors are the most discussed methods for exploiting IoT devices. To determine how easily these vulnerabilities can be exploited, I used the IoT search engine, Shodan.
4.3.2 Shodan: Discovering Vulnerable IoT Devices
Using Shodan, I followed these steps for my analysis:
@@ 4.3.2 Shodan: Discovering Vulnerable IoT Devices
Figure 3: Example of vulnerable temperature-sensing device found through Shodan
Figure 4: Example of vulnerable sensor devices found through Shodan
Figure 5: Example of vulnerable pressure-sensor device found through Shodan
Figure 6: Example of full access admin control panel of pressure-sensing device found through Shodan
Upon executing the Python script detailed in Listing 1, I obtained 582 results for sensors. In the try block of the code, I captured the number of results and filtered information such as company name, IP address, port, operating system, and data details, including content type, authentication protocol, and location specifics like country and city. The results from my Shodan search provided comprehensive information about the devices, including company name, location, IP address, and port. This level of detail facilitates various forms of analysis. First, I filtered the results to include only those with port 8080 and directly accessed some of the IP addresses. As shown in the Figure 3, I encountered a page requesting login credentials. This is concerning because it exposes the device to a range of security vulnerabilities, such as cross-site scripting, denial-of-service attacks, SQL injection, unexpected application termination, and arbitrary code execution. Additionally, the connection was not secured over SSL. Second, I manually examined the IP addresses providing direct access. As illustrated in Figure 4 and Figure 5, I gained access to a pressure-sensing device named ’Pbar-box4’ and a temperature-sensing device named ’Byt.’ I had full control over these devices, allowing me to view past temperature readings and manipulate administrative settings, including changing email addresses where data was being sent. Figure 6 demonstrates an example of the control panel for the ’Byt’ device.
5 Conclusions
Analyzing and collecting Dark Web information is exceedingly challenging due to cyber-attackers’ ability to conceal their identities and erase traces of their activities online. The sheer volume of web information complicates the task of obtaining a comprehensive understanding of their operations, making it even more difficult to predict their next targets. In this article, I propose a methodology to address these challenges. By employing web mining, content analysis, and visualization techniques, the methodology leverages various information sources and analyzes 23 hacker websites for IoT vulnerabilities. Information clustering is used to differentiate between novice and experienced hackers on the web. My detailed evaluation indicates that the methodology produced promising results, which could significantly aid in discovering new vulnerabilities and their sources.
For IoT, I propose the following solutions:

### v4

@@ Abstract
Index Terms:
Internet of Things (IoT), Dark Web, IoT Vulnerability, Machine Learning in Cybersecurity, Threat Intelligence.
I Introduction
@@
In this paper, we focus on the Internet of Things (IoT) as a vulnerability due to its expanding presence and susceptibility to exploitation [2]. By employing the structured methodology outlined in this study, we evaluate the degree to which cybercriminal groups target IoT systems. Through our analysis of 23 criminal websites on the Dark Web, we discovered that even seemingly harmless devices, such as small sensors, can cause substantial damage when compromised. IoT opens up an entirely new realm for cybercriminals to exploit. These hacker groups generally do not discriminate in their targets—they exploit any weakness or vulnerability they can find.
@@
The remainder of the paper is organized as follows: Section II provides an overview of the concepts introduced in this paper; Section III outlines a methodology for collecting and analyzing Dark Web data, and identifying vulnerabilities; Section IV details how this methodology was applied in a case study of IoT using Shodan, the IoT search engine; and the final section concludes the study and explores potential solutions.
II Literature Review
II-A Dark Web Utilization
@@
Hackers, criminals, and terrorists leverage the web to facilitate their illicit activities. Since the emergence of the Dark Web, numerous new marketplaces have surfaced, offering a wide and often alarming array of products, including prescription pills, meth, heroin, speed, crack, guns, stolen identities, gold, and erotica. The Dark Web also serves as a platform for spreading extremism, facilitating communication among terrorists and criminals, and sharing knowledge about illegal hacks and attacks, thereby enabling harmful activities that threaten society. Most of these markets operate using Bitcoin and other cryptocurrencies, which allow for anonymous online transactions. According to one study, criminals and terrorists use Bitcoin to sell drugs and subsequently purchase guns and ammunition. During our exploration of the dark market known as Hansa Market, we identified a website called Arms and Ammunition, which poses significant threats to nations both internally and externally. This underscores the importance of developing robust threat intelligence.
II-B Discovering IoT Devices Through Shodan
We focused on IoT devices due to their growing integration into daily life and critical infrastructure. Like mobile phones, these devices are increasingly capable of learning and predicting user behavior. Despite their utility, many IoT devices operate on low-power processors and custom operating systems that often lack robust encryption capabilities, making them inherently vulnerable.
For example, a smart thermostat not only regulates temperature but also stores behavioral data on centralized servers. While a compromised thermostat in a home may seem low-risk, similar vulnerabilities in industrial control systems can lead to serious disruptions. Attackers exploiting such weaknesses could manipulate systems or target individuals based on behavioral patterns, posing threats to both personal safety and public infrastructure.
To identify exposed IoT devices, we used Shodan, a search engine that scans the internet for connected devices with open ports. Shodan can discover a wide range of systems, from webcams and servers to industrial control devices and power grids. Its effectiveness has been demonstrated in prior research, such as a study in which newly connected devices were indexed by Shodan within 14 days, revealing their IPs, ports, and system metadata [1].
Shodan gathers information through service banner interrogation—extracting data provided by devices in response to connection requests. With powerful filters such as IP address, port, location, and operating system, Shodan offers both attackers and defenders a valuable tool for identifying vulnerable systems and enhancing threat intelligence.
II-C Developing Cyber Threat Intelligence
@@
Our research focused on determining whether the systematic collection and processing of Dark Web content could support the development of actionable threat intelligence to protect vulnerable IoT devices. We examined how insights from underground hacker forums, marketplaces, and other illicit platforms could enhance understanding of emerging cyber threats targeting IoT ecosystems. The remainder of this paper details the methodologies we employed for data collection, the techniques used to filter and refine relevant information, and the analytical approaches applied to transform raw data into meaningful threat intelligence. By outlining these processes, we demonstrate how Dark Web intelligence can be leveraged to strengthen IoT security measures and mitigate cyber risks before they materialize into active threats.
III A Methodology for Collecting and Analyzing Dark Web Information
III-A The Methodology
@@
To address the diverse threats posed by the information sources used by hackers, criminals, and terrorists, we propose a comprehensive threat intelligence framework. This framework integrates Dark Web data aggregation, machine learning-driven anomaly detection of IoT vulnerability patterns, and human-AI collaboration for real-time cyber-attack prediction.
@@
Unlike prior approaches [25], our method correlates hacker discussions (e.g., IoT exploitation trends) with Shodan-derived device vulnerabilities to prioritize high-risk targets. Figure 1 illustrates this methodology, designed to assist investigators in obtaining Dark Web intelligence through diverse sources, collection techniques, filtering, and analysis [14].
  • Information Sources cover a wide range of hacker-related platforms, from public search engines to sites maintained by hacker groups, hacktivists, terrorists, or supporters—often requiring domain expertise for access.
  • Collection employs systematic crawling starting from seed URLs or keywords, leveraging various web scraping tools to gather relevant IoT threat data.
  • Back-link Search uses search engines like Google to trace web pages linking to hacker domains, helping uncover supporters who glorify or coordinate attacks.
  • Personal/People Data Search through portals such as Yahoo! and MSN reveals profiles with links that expose hidden connections among hacker communities.
  • Meta-searching aggregates results from multiple search engines using related keywords to refine and prioritize relevant information about groups like ”Anonymous.”
  • Filtering requires domain and linguistic expertise to remove irrelevant data, considering the history, language, and symbolism of cyber-attackers.
  • Analysis aids investigators by profiling IoT exploit motives, clustering hacker entities, mapping risk topographies, scoring threats based on hacker interest and exposure, and discovering vulnerabilities linked to specific targets.
III-B Application of the Methodology
@@
As discussed earlier, our proposed methodology integrates a combination of advanced data mining, web mining, and machine learning techniques while leveraging human domain expertise to refine their application. This hybrid approach enhances the ability to systematically identify and monitor criminal groups, analyze their discussions, and extract intelligence relevant to emerging cyber threats. By incorporating both automated data processing and human analytical oversight, our methodology balances computational efficiency with contextual accuracy, enabling more precise intelligence gathering. Ultimately, this semi-automated framework presents a scalable solution for uncovering hidden threat actors and understanding the evolving landscape of cybercrime on the Dark Web.
IV Case Study: IoT Scanning to Predict Cyber Attacks
To illustrate the effectiveness and applicability of the methodology, we applied it to the collection and analysis of web-based activities by malicious actors targeting IoT devices. These devices are especially vulnerable and projected to number 26 billion by 2030. We describe our approach in Figure 1.
IV-A Collection
@@
To gather data, we first identified several suspicious URLs using web searches, referencing dark web markets. We conducted web spidering using the ’Open Web Spidering’ tool. For instance, starting with keywords like “freedom fighters” and “digital Robin Hood,” we found four intriguing URLs: HackHound (a forum for hackers), Du Bist Anonymous (a website supporting the Anonymous group), CiHad.net (a suspicious website associated with a terror group), and Intel Exchange (a political ideology website). Using these categories as references, we uncovered additional websites during the data collection process.
We then performed the same searches on dark web search engines such as Torch, Grams, and Ahmia. Although the results were scattered, we stored them nonetheless, as they primarily pointed to dark web markets. Manually browsing these markets revealed offers of both hacking services and services required for hacking. We selected keywords such as “hacking services,” “free world,” “personal army,” and “Internet of Things,” which yielded a variety of relevant results. Overall, we collected information from hundreds of sources, and a promising future direction would be to study how different data collection techniques affect the quality and depth of subsequent analysis.
IV-B Filtering
We conducted two rounds of filtering to refine the dataset. In the first round, we manually excluded websites that were clearly unrelated to cybersecurity or hacking, such as news portals, gaming platforms, and adult content sites. We retained websites hosting forums, technical blogs, ideological discussions, and dark market listings offering hacking tools, data leaks, or illicit services such as ammunition. This process reduced the dataset to thirty-two websites.
In the second round, we refined the selection further. Given limitations in language coverage and domain-specific nuance, we manually removed sites that appeared redundant, linguistically inaccessible, or thematically distant from our focus. To support this process, we applied natural language processing (NLP) and basic machine learning (ML) techniques to tag security-relevant keywords and compute a threat relevance score for each website. This score was generated using TF-IDF-based feature extraction and a logistic regression model trained on a small labeled dataset comprising known malicious and benign cybersecurity forums. Websites with higher scores—indicating stronger alignment with cyber threat indicators—were prioritized for deeper analysis. This hybrid manual-automated filtering approach helped reduce subjectivity and improved consistency in identifying high-risk content.
IV-C Analysis
IV-C1 Classification and Usage of Websites
@@
We conducted two rounds of filtering. First, we manually removed unrelated sites, such as news, games, and porn sites, which had no reference to hacking. We retained websites of forums, blogs, ideological discussions, and dark markets for hacking and ammunition services. After this round, thirty-two sites remained.
Second, given our limited domain expertise and knowledge of languages, we initially manually removed some websites that felt less relevant and were similar to English forum websites in our results. Subsequently, we used natural language processing (NLP) and machine learning (ML) for keyword tagging and threat scoring.
@@
We performed clustering and classification of the remaining 23 websites. These websites provided a diverse range of information necessary for our analysis. Table I lists these websites with hyperlinked URLs and descriptions.
@@
TABLE I: Websites collected with the implementation of the methodology.
We followed the following steps to analyze “IoT Hacking” or “hacking IoT devices” discussions on some of these websites:
  • We filtered cybersecurity hacking discussion forums from the list above, ignoring the rest of the websites.
  • We manually browsed through these forums and searched for the keywords listed.
  • We gathered statistics as recent as possible (up to 2024) and generated a graph (Figure 2) and a table (Table II).
The graph in Figure 2 displays the percentage of IoT hacking-related discussions occurring in the seven forums we selected for analysis. The x-axis shows the names of these forums, while the y-axis shows the percentage of IoT hacking discussions. The ”Hacker Web” forum, which aggregates 18 major forums, has 3.3% of its posts discussing IoT hacking. Although this percentage is lower, it does not imply fewer discussions, given the forum’s size. The other forums show 12% to 30% of their topics discussing IoT hacking.
@@
TABLE II: This table shows the keywords searched and data found on the mentioned websites.
By manually browsing through these forums, we analyzed the most frequently discussed topics related to IoT hacking. We selected four websites listed in Table II, searched for the keywords in the ’Searched’ column, and recorded the results and total post counts in the ’Found’ and ’# of Posts’ columns, respectively. Our analysis revealed that botnets, malware, and sensors are the most discussed methods for exploiting IoT devices. To determine how easily these vulnerabilities can be exploited, we used the IoT search engine, Shodan.
IV-C2 Shodan: Discovering Vulnerable IoT Devices
Using Shodan, we followed these steps for our analysis:
Upon executing the Python script detailed in Listing 1, we obtained 582 results for sensors. In the try block of the code, we captured the number of results and filtered information such as company name, IP address, port, operating system, and data details, including content type, authentication protocol, and location specifics like country and city. The results from our Shodan search provided comprehensive information about the devices, including company name, location, IP address, and port. This level of detail facilitates various forms of analysis.
@@ IV-C2 Shodan: Discovering Vulnerable IoT Devices
(a) Temperature-sensing device
(b) Other sensor devices
(c) Pressure-sensor device
(d) Admin control panel
Figure 3: Examples of vulnerable IoT devices discovered via Shodan
First, we filtered the results to include only those with port 8080 and directly accessed some of the IP addresses. As shown in Figure 3(a), we encountered a page requesting login credentials. This is concerning because it exposes the device to a range of security vulnerabilities, such as cross-site scripting, denial-of-service attacks, SQL injection, unexpected application termination, and arbitrary code execution. Additionally, the connection was not secured over SSL.
Second, we manually examined the IP addresses providing direct access. As illustrated in Figures 3(b) and 3(c), we gained access to a pressure-sensing device named ’Pbar-box4’ and a temperature-sensing device named ’Byt.’ We had full control over these devices, allowing us to view past temperature readings and manipulate administrative settings, including changing email addresses where data was being sent. Figure 3(d) demonstrates an example of the control panel for the ’Byt’ device.
V Conclusions
Analyzing and collecting Dark Web information is exceedingly challenging due to cyber-attackers’ ability to conceal their identities and erase traces of their activities online. The sheer volume of web information complicates the task of obtaining a comprehensive understanding of their operations, making it even more difficult to predict their next targets. In this article, we propose a methodology to address these challenges. By employing web mining, content analysis, and visualization techniques, the methodology leverages various information sources and analyzes 23 hacker websites for IoT vulnerabilities. Information clustering is used to differentiate between novice and experienced hackers on the web. Our detailed evaluation indicates that the methodology produced promising results, which could significantly aid in discovering new vulnerabilities and their sources.
For IoT, we propose the following solutions: